find the area of the space enclosed by two concentric circles of radii 15cm & 13cm.
Answer by edjones(8007)

pi*r^2=A
pi*15^2=225*pi
pi*13^2=169*pi
-------------subtract
=========56*pi cm Ans
.
Ed